OpenStudy (anonymous):

PLEASE HELP! I don't understand this at all. A series RLC circuit is connected with source providing amplitude voltage of V= 120 V. The amplitude voltages measured accross the capacitor and inductor were found to be V(c)=20 V and V(L)=80 V repectively. The resistance used has R=150 ohms. a) what is the average power supplied by the source?

OpenStudy (sidsiddhartha):

what is the frequency of the suppy, is it 50hz? without it we can have some difficulties

OpenStudy (anonymous):

It doesn't say, thus one of the reasons I am having problems

OpenStudy (sidsiddhartha):

ok then u have to assume it because without it we cannot find values of C and L as \[X_c=1/\omega_c=\frac{ 1 }{ 2 \pi f }\\and\\X_l=2 \pi f*L\]

OpenStudy (sidsiddhartha):

it will be \[X_c=1/2 \pi f C\]

OpenStudy (sidsiddhartha):

ok lets say f=50 hz now can u do it?

OpenStudy (sidsiddhartha):

 thats the situation